At the park there is a pool shaped like a circle. A ring shaped path goes around the pool. It's inner diameter is 20 yd and outer is 28 yd. If one gallon of coating can cover 7 yd squared, how many gallons of coating do you need?
Answer by nerdybill(7384)

At the park there is a pool shaped like a circle. A ring shaped path goes around the pool. It's inner diameter is 20 yd and outer is 28 yd. If one gallon of coating can cover 7 yd squared, how many gallons of coating do you need?
"area that needs painting" = "area of outside ring" - "area of inside ring"
"area that needs painting" = (pi)(28)^2 - (pi)(20)^2
"area that needs painting" = (pi)(28^2 - 20^2)
"area that needs painting" = (3.14)(28^2 - 20^2)
"area that needs painting" = (3.14)(384)
"area that needs painting" = 1205.76 square yards
.
paint needed:
1205.76/7 = 172.25
or
173 gallons of paint